OpenStudy (anonymous):

would (3/4)-(-2/3)= 0.5 or negative -0.5

OpenStudy (anonymous):

its 17/12

OpenStudy (anonymous):

1/5/12

OpenStudy (anonymous):

Wait what O_o

OpenStudy (anonymous):

17/12

OpenStudy (anonymous):

3/4 - 2/3 = (3*3)/12+ (2*4)/12 = 17/12